OpenStudy (anonymous):

Solve. 12x^2 + 7x = 10

OpenStudy (anonymous):

take out an x to get x(12x+7)=10. Then split into 2 equations to get x=10 and 12x+7=10. Then solve for x. So 12x+7=10 move the 7 over through subtraction and you get 12x=3. Then just divide 3 by 12 to get 1/4. So x = 10 and x =1/4

OpenStudy (anonymous):

\[12x ^{2} + 7x - 10 =0\] factoring (3x-2) (4X+5) =0 3X=2 4X = -5 X = 2/3 X = -5/4

OpenStudy (anonymous):

ignore mine.

OpenStudy (anonymous):

@abigail you cant split equations if they are not equal to 0 ^^

OpenStudy (anonymous):

ya i realized it after i put it up. lol

OpenStudy (anonymous):

@pineda100 can you check another question i asked that was answered?